OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 15 and a13= -57

OpenStudy (anonymous):

oh gosh letme think

OpenStudy (anonymous):

what grade are you in :O

OpenStudy (anonymous):

9

OpenStudy (anonymous):

oh im in 7th, thats why i dont get it v.v sorry

OpenStudy (caominhim):

a1 = 15 a13 = -57 in 12 iterations you dropped (15+57) = 72 therefore you dropped 72/12 = 6 per iteration \[A _{n} = A _{n-1}-6\] therefore \[A _{32} = 15-6(31)\]

OpenStudy (anonymous):

thank you